Which is shorter, a 95% z-interval for the mean or a 95% t-interval for the mean? Is one of these always shorter, or does the outcome depend on the sample?


1
Expert's answer
2021-01-18T09:05:00-0500

Confidence Intervals for "\mu"

For "n\geq30"


"\bar{X}\pm z\times\dfrac{s}{\sqrt{n}}"

Use the Z table for the standard normal distribution.


For "n<30"


"\bar{X}\pm t\times\dfrac{s}{\sqrt{n}}"

Use the t table with "df=n-1"

"\alpha=0.05"

"z=1.96"

"df=1, t=12.706"

"df=10, t=2.228"

"df=20, t=2.086"

"df=30, t=2.042"

t-value is greater than z-value.

A 95% z-interval for the mean is always shorter than  a 95% t-interval for the mean.